OpenStudy (cicilybailey14):

I medal!!!!! A direct variation includes the point (-8, 2). Which is an equation of the direct variation? y = -4x y = x/-4 2y = -8x -8y = x + 2

OpenStudy (danjs):

Direct variation means this form: y = k*x Use the point (x,y) = (-8,2) in that, and figure out the value of k

OpenStudy (danjs):

2 = k * (-8) k = 2/(-8) = -1/4

OpenStudy (danjs):

so the direct variation is y = k*x y = (-1/4) *x

OpenStudy (danjs):

or in the form of the answers \[y = \frac{ x }{ -4 }\]

OpenStudy (cicilybailey14):

Thanks! :")